# Turnstile area as a measure for chaotic transport in magnetic confinement fusion devices

### v1

@@ II.3 Maps defined by three-dimensional fields
In such a stochastic region, trajectories explore large portions of $\psi$, as is seen from the mixing of different colored trajectories. This can have the effect of mixing between hot and cold regions in the plasma which is detrimental to confinement. Even though there is mixing, we still see the gradient of color from red to blue in the stochastic region, and one can ask the question which fraction of the field lines near the plasma will reach regions further out, and how many iterations of the map this will on average take.
Such questions are important for designing a fusion reactor, because the hot plasma is confined by the magnetic field, and heat and particle transport along the field lines can be up to 10 orders of magnitude faster along the field, the trajectories of the system than perpendicular to them. If we place a wall around the plasma, where do the trajectories intersect the wall, on average how far along the trajectory (how many iterations of the map) does it take for the plasma to reach the wall- these are questions that require knowledge of the magnetic topology.
@@
For this, we first calculate the fixed points of each of the island chains. For the period 3 island chain, we use the Newton’s method in equation (7) applied to the 3-map. We find an elliptic period-3 orbit in the center of the island, shown with a blue dot in figure 2, and a hyperbolic period-3 orbit in between marked with a blue x. Similarly we find the x-points in the period 5 and period 2 orbits.
@@
We calculate the manifolds of all hyperbolic points. First we choose one of the $n$ points of the periodic orbit, and evaluate the linearized mapping $\mathsf{M}$ and it’s eigenvectors. We choose point a distance of $10^{-5}$ (using the Euclidean norm on $\psi,\vartheta$) from the fixed point along the unstable eigenvector, and evaluate the map of that point. After confirming that this is in the linear regime, we choose 50 points logarithmically spaced between the initial point and the mapped to point. We then calculate the trajectories of these 50 distinct points for 12 applications of the $n$-map (12 was found to be a sufficient number to illustrate the manifold). If $\mathbf{v}$ is an unstable eigenvector of $\mathsf{M}$, then $-\mathbf{v}$ is as well, so we calculate the second unstable manifold as well. Since the first point of this segment of points maps to the last, the result is a continuous line, that starts straight but becomes more convoluted upon higher iterates of the map. The result of this calculation are the unstable manifolds of of the x-points, shown in magenta.
@@
Figure 3: Illustration of transport through the barrier created by the first period 3 island chain in the Tokamap. Trajectories enter the zone around the island through a ’turnstile’ created by the intersecting manifolds on the inner heteroclinic connection. The orange set originates from the plasma region (yellow areas), and is mapped forwards into the region around the island (red areas). There is an equivalent set that is mapped from the island zone into the plasma zone. The transport away from the island is governed by another turnstile illustrated with the cyan area, whose backwards map is shown in green, and forwards map in blue.
In order to understand the chaotic transport of field lines through an island chain, we take a zoomed-in look at a single island indicated by the box in figure 2. Figure 3 shows the two period-3 fixed points labeled $\mathbf{x}^{1}_{0}$ and $\mathbf{x}^{2}_{0}$ and the stable and unstable manifolds associated with them. The area around the island o-point, and bounded by the manifolds, is called the island resonance zone (for the exact definition see [37]). Consider the orange area labeled $\Sigma_{\text{inner}}$, on the bottom, defined by the intersections of the inner manifolds. If we apply the forwards map $f^{3n}(\Sigma_{\text{inner}})$, it is mapped to the to the red areas, and when we apply the backwards map, it is mapped to the yellow areas.
@@
Locating the heteroclinic orbits can be done by computing the complete manifolds, and locating their intersections. Calculating the manifolds is a computationally costly endeavor, as it requires integrating a large number of trajectories for many iterations. For iterated maps this is not an issue, but if the map is calculated by the numerical integration of trajectories, the computational cost can be significant. Therefore, we implement an algorithm to find these heteroclinic orbits directly with many fewer required integrations.
@@
(here the superscript denotes the fixed point label). We choose $\epsilon$ small enough to be in the linear regime, here $10^{-6}$, and we then compute $f^{mn_{\rm map}}(\mathbf{x}_{u})$, where $m=3$ is the order of the fixed point, and $n_{\rm map}$ is the number of times this initial point is mapped back to the same intersection.
@@
and calculate $f^{-m\cdot n_{\rm map}}(\mathbf{x}_{s})$.
If $\mathbf{x}_{s}$ and $\mathbf{x}_{u}$ are heteroclinic points, then $f^{m\cdot n_{\rm map}}(\mathbf{x}_{u})=f^{-m\cdot n_{\rm map}}(\mathbf{x}_{u})$, but this is unlikely to be the case. We therefore minimize the components of:
$\Delta=f^{m\cdot n_{\rm map}}(\mathbf{x}_{0}^{1}+\epsilon_{u}\cdot\mathbf{v}_{u}^{1})-f^{-m\cdot n_{\rm map}}(\mathbf{x}_{0}^{2}+\epsilon_{s}\cdot\mathbf{v}_{s}^{2}).$ (16)
@@
We choose $n_{\rm map}=8$ for both the stable and the unstable manifold integration, and the values of $\epsilon_{s,u}$ are initially set to $10^{-6}$. The root-finding converges resulting in a heteroclinic trajectory with 16 points. These are shown in the top panel of 2.
The heteroclinic trajectories thus found are shown in figure 3 as the blue circles (squares) for the minimizing (minimax) orbits respectively.
@@ V Chaotic transport in a perturbed tokamak
In this section, we will analyze chaotic transport in a tokamak-like field that has been perturbed. We will first describe the analytical expressions used to calculate an axisymmetric, tokamak-like field and analytical perturbations to it. We calculate the manifolds in this perturbed system and describe an algorithm to find the homoclinic orbits, the intersections of these manifolds Using the homoclinic orbits, we describe an efficient algorithm to calculate the turnstile area using an action principle by MacKay, Meiss and Persival [16].
In a three-dimensional field, the fast-evaluating function such as (9), is replaced by an expensive integration of the trajectory as in (1). Though the variables $(\psi,\vartheta)$ are replaced with $(R,Z)$, all of the discussed algorithms remain the same. We will start with a axisymmetric field, that resembles the configuration in a Tokamak. Calculation of the manifolds in tokamak geometry has been studied theoretically by Punjabi and Boozer [3], Wei and Liang [38], and their signatures have been shown in experiments [39]. Here we build upon this work by implementing our efficient method of calculating the turnstile area, and hence a measure for the degree of stochasticity and stochastic transport.
@@
which is chosen such as to give a simple expression for the rotational transform $i$ or it’s inverse, $q=1/\imath$ which is more often used in tokamak physics:
@@
Here $K$ and $E$ the complete elliptic integral of the first and second kind respectively, $\alpha^{2}=(R_{l}-R)^{2}+(Z-Z_{l})^{2}$, $\beta^{2}=(R_{l}+R)^{2}+(Z-Z_{l})^{2}$, and $k=1-\alpha^{2}/\beta^{2}$.
Equations (17) and (21) are implemented numerically in terms of their respective vector potentials, and the magnetic field is calculated through $\mathbf{B}=\nabla\times\mathbf{A}$ using automatic differentiation provided by JAX [41]. We choose $R_{0}=6$, $q_{a}=1.16$, $s=1.2$ for the axisymmetric field, and for the separatrix coil we choose $R_{l}=6$, $Z_{l}=-5.5$, $I=10\pi/\mu_{0}$. The resultant field is integrated numerically, and a Poincaré section is shown in figure 4 a). The fixed points of the 1-map of this field are also located, and these are the magnetic axis shown with a blue dot, the location of the separatrix coil shown with a red square, and a hyperbolic x-point in between, shown with an orange cross.
Figure 4: Poincaré section of the axisymmetric analytical field and the result of applying a perturbation. a) Axisymmetric field $\mathbf{B}_{\rm axi}$ with the o-point of the main plasma in blue, the separatrix coil location indicated in red, and the x-point it creates identified by an orange cross. The unstable and stable manifolds that enclose the plasma overlap, and are shown by a magenta and blue line respectively. b) Effect of applying a $m=6$, $n=1$ perturbation field. The direction the perturbation field is indicated by arrows located on the unperturbed separatrix and on a circle in the plasma region. The perturbation creates islands inside the plasma region, and stochastization around the separatrix. The stable and unstable manifolds (magenta and blue respectively) no longer overlap. c) safety factor $q$ (inverse of rotational transform $i$ measured along a horizontal line from the axis, in both the unperturbed (blue) and perturbed field (orange). The islands in the perturbed field correspond with the regions of flat $q$ at integer values of $q$.
Because the fields given by equations (17) and (21) are axisymmetric, the resultant Poincaré section in fig. 4 a) does not exhibit any chaos, and there are no turnstiles. Now we add a perturbation that depends on the toroidal angle $\phi$ with the form:
@@
where $\phi$ is the usual toroidal angle and $\theta$ is the poloidal angle; $\tan(\theta)=(Z-Z_{0})/(R-R_{0})$ (with $(R_{0},Z_{0})$ the location of the magnetic axis, blue circle in fig. 4 a)). The radial distribution $\mathfrak{f}(\rho)$ of the perturbation can have any form, we choose one that places the perturbation smoothly everywhere in the configuration by using a Maxwell-Boltzmann probability density function with distribution parameter $d$:
$\mathfrak{f}(\rho,d)=\frac{\sqrt{2}}{\sqrt{\pi}}\frac{\rho^{2}}{d^{3}}\exp{\frac{-\rho^{2}}{2d^{2}}}$ (23)
@@
$\mathbf{B}_{\rm pert}$ lies in the $R,Z$-plane, and is tangent to contours of constant $R^{2}A^{\phi}_{\rm pert}$. The perturbation field oscillates $m$ times in one $R,Z$-plane, and this pattern rotates $n$ times as the angle goes from 0 to $2\pi$. We choose $\phi_{0}=0$, $\theta_{0}=0$, distribution width $d=2$, and mode numbers $m=6$ and $n=1$. This is shown in figure 4.
When this perturbation is applied, the field becomes stochastic, as is seen in figure 4 (right). Islands have opened up on several magnetic surfaces, and the region around the separatrix surface has become chaotic. The x-point is no longer clearly visible in the Poincaré section. Nevertheless, the fixed point is still there, and it is found using the Newton method described in section III.1, and it is indicated with a orange cross.
In section IV we considered a heteroclinic intersection, where stable manifolds of one fixed point intersect with the unstable manifolds of a second point in the same orbit, but here, we can consider the simpler where the stable and unstable manifolds originate from the same point; a homoclinic connection. We find the homoclinic trajectory in this integrated map using the same algorithm as described in IV.1, and calculate the manifolds in the same way as described in section IV. The results of this calculation are shown in figure 4 b).
@@ V.1 Turnstile area calculation
Macay, Meiss, and Percival [16] describe a simpler method to calculate the turnstile area. Here we give a brief overview of that method, but for a full description and definitions of terms, we refer the reader to the review by Meiss [37] Their method is based on the Lagrangian form $\lambda$ calculated on homoclinic orbits. They show that if the map is exact, $\lambda=F(x,x^{\prime})$ (where $x$ is a point in the section and $x^{\prime}$ is the point mapped to) which is one of the standard generating functions for canonical transformations. They go on to show, for a homo (or hetero)clinic connection, which has two homo (hetero)clinic trajectories, $m_{t}$ the minimax orbit and $h_{t}$ the minimizing orbit, that the homoclinic lobe area can be calculated directly from:
@@
In the case of magnetic fields, the Lagrangian form is calculated by integrating the vector potential along the field line [THIS IS NOT EXACTLY RIGHT]:
$\lambda(m_{t})=\int_{0}^{mT}\mathbf{A}(\mathbf{x})\cdot\mathrm{d}l.$ (26)
where the integration is started at $\mathbf{x}=m_{t-1}$ and the integration is performed along the field line.
It seems rather remarkable that the computation of a complicated area in phase space can be performed by the simple summation (equation (25)) of integrals (equation (26)). Though the formulation in terms of Lagrangian forms and area forms holds in any dimension, in the case of a three-dimensional flow and a two-dimensional map, this formula has an intuitive explanation.
Figure 5: Illustration of the calculation of a turnstile area. The area $\Sigma_{n}$ shown in gray is bounded by turnstile sections $\gamma_{n}$ and $\tau_{n}$ between homoclinic points, is given by the integral of the vector potential over the curve $\tau_{n}\cup\gamma_{n}$. This integral can be modified to run over $\tau_{n}\cup\xi_{n-1}\cup\gamma_{n-1}\cup\zeta_{n}-1$. This can be repeated until $\gamma_{n-i}$ is in the linear regime of the fixed point. Similarly the integration over $\tau_{n}$ can be moved forward to $\tau_{n+j}$ in the linear regime of the slow manifold. (bottom): illustration of the turnstile area calculation where the intersecting manifolds have been laid flat.
Figure 5 illustrates how the calculation of the turnstile lobe can be understood. To calculate the flux through the dark shaded area, which without loss of generality we label the $n$-th occurrence of the lobe, one can use Stokes’ theorem and integrate the vector potential $A$ over $\gamma_{n}$, a section of the unstable manifold, and $\tau_{n}$, a section of the stable manifold:
@@
We can now shift the integral over the unstable manifold by transporting it along the field to the next intersection. This integral now becomes:
@@
Instead of integrating the vector potential over two curve segments, whose geometry has to be found by integrating many trajectories, we can instead identify two specific trajectories along which to integrate the vector potential. One issue in the practical application of equation (26) is when to terminate the sum. In practice only about $N=$10-20 applications of the map suffice to bring a point from the linear regime of the unstable manifold into the linear regime of the stable manifold. Increasing the number of points in the homoclinic trajectories will make the sum converge, but at a large computational cost. Instead, when we are in the linear regime, we can close the sum by evaluating $\int\mathbf{A}\cdot\mathrm{d}l$ along the straight line connecting the last homoclinic points.
@@
Figure 6: The effect of perturbations on the turnstile area using the same axisymmetric field as in figure 4. a-l: Each column has a different poloidal mode number $m$, and each row has a different amplitude of the perturbation. Different mode numbers affect the field and chaotic transport differently. Low $m$ mode numbers are resonant with the surfaces inside, and create stochastic fields throughout the volume. m: Turnstile area vs perturbation amplitude. The response to perturbation is mostly linear, but can act nonlinearly at very high amplitudes. n: Turnstile area at a perturbation amplitude of 0.004. Certain modes are more efficient at increasing turnstile area and generating stochastic transport than others.
@@
We now use this algorithm to compare the effects of varying the perturbation mode numbers and perturbation amplitude on the turnstile area in the tokamak toy model equilibrium. To compare different perturbations, we normalize the perturbation strength the axisymmetric field strength, by evaluating the maximum of $|B_{\rm pert}|/|B_{\rm axi}|$ on the unperturbed separatrix. Poincaré sections at three amplitudes, for the mode numbers $m=2,4,7,9$ ($n$ always 1) are shown in figure 6 a). All perturbations create chaotic regions, but the perturbations with lower $m$ create more stochasticity in the core. The rotational transform crosses the rational values $m/1$ for those $m<7$ and the mode couples with resonant surfaces there.
@@
The size of the turnstile area as a function of the mode number is shown in figure 6 b, The area scales approximately linearly with the applied perturbation, though some modes show nonlinearity at high amplitudes. The method is able to calculate turnstile areas over a large range of perturbation amplitudes, and at the maximum applied perturbations the perturbation field is is a significant fraction of the unperturbed field strength. We thus have a reliable method to evaluate the turnstile area over a broad range of perturbation amplitudes, from nearly integrable fields, to moderately chaotic fields up to almost completely stochastic fields.
@@
Different perturbations are more or less effective at creating stochastic transport on the separatrix. To compare this, we plot the turnstile area at a fixed perturbation amplitude of $\max\left(\tfrac{|B_{\rm pert}|}{|B_{\rm axi}|}\right)=0.0040$ for all different modes. This is shown in figure 6 n.
@@
Understanding which modes are most effective at creating stochasticity in fields is an important and long-standing area of research in tokamak physics, where so-called Resonant Magnetic Perturbation (RMP) [42] coils are placed around the vessel to create a stochastic edge. Computing the lobe structures is key to predicting the heat footprints on divertor targets [43]. We expect that the turnstile area calculation could be an effective tool in designing RMP fields. Given the potential risk posed by runaway electrons for future experiments, there is a big effort in studying mitigation methods, and one proposal are Passive Runaway Electron Mitigation Coil (PREMCs) [44]. Here a shaped coil is placed near the vessel, and the current induced in it by the abrupt changes in flux brought on by the disruptive events preceding the formation of a dangerous beam of near-relativistic electrons, will induce a current in the coil which stochasticizes the field, disrupting the beam. Such a coil could be optimized to create maximum stochasticity at desired locations.

### v3

@@ II.3 Maps defined by three-dimensional fields
In such a chaotic region, trajectories explore large portions of $\psi$, as is seen from the mixing of different colored trajectories. This can have the effect of mixing between hot (small $\psi$) and cold (large $\psi$) regions in the plasma which is detrimental to confinement. Even though there is mixing, we still see the gradient of color from red to blue in the chaotic region, indicating that only a fraction of the red field lines near the plasma will mix in the blue regions further out.
The amount of field lines that is exchanged between the region on one side of an island chain and the other side is important, because heat and particles are transported along these trajectories. Where do the trajectories that pass through the system end up on the wall?
@@
To analyze this transport we first find the hyperbolic fixed points of each of the island chains. For the period 3 island chain, we use the Newton’s method in equation (7) applied to the 3-map, with initial guess chosen by eye from the Poincaré map. The three points on the period-3 hyperbolic trajectory are marked with a black ’x’ in figure 2. Similarly we find and mark the hyperbolic fixed points in the period 5 and period 2 orbits.
@@
We calculate the manifolds of all hyperbolic points. First we choose one of the $n$ points of the periodic orbit, and evaluate the linearized mapping $\mathsf{M}^{n}$ and its eigenvectors. We choose point a distance of $10^{-5}$ (using the Euclidean norm on $\psi,\vartheta$) from the fixed point along the unstable eigenvector, and evaluate the map of that point. After confirming that this is in the linear regime, we choose 50 points logarithmically spaced between the initial point and the mapped to point. We then calculate the trajectories of these 50 distinct points for 12 applications of the $n$-map (12 was found to be a sufficient number to illustrate the manifold). If $\mathbf{v}$ is an unstable eigenvector of $\mathsf{M}$, then $-\mathbf{v}$ is as well, so we calculate the second unstable manifold as well. Since the first point of this segment of points maps to the last, the result is a continuous line, that starts straight but becomes more convoluted upon higher iterates of the map. The result of this calculation are the unstable manifolds of of the x-points, shown in magenta.
@@
Figure 3: Illustration of transport through the barrier created by the first period 3 island chain in the Tokamap. Trajectories enter the zone around the island through a ’turnstile’ created by the intersecting manifolds on the inner heteroclinic connection. The orange set labeled $\Sigma_{\text{inner}}$ originates from the plasma region (yellow areas), and is mapped forwards into the region around the island (red areas). There is an equivalent set that is mapped from the island zone into the plasma zone. The two fixed points from whence the manifolds originate are labeled $\mathbf{x}_{0}^{1}$ and $\mathbf{x}_{0}^{2}$, and their two eigenvectors that are used to calculate the inner heteroclinic orbits are shown by the red (blue) arrows labeled $\mathbf{v}_{u}^{1}$ ($\mathbf{v}_{s}^{2}$) respectively. The heteroclinic orbits on the intersections of the manifolds are shown with blue dots. The transport away from the island is governed by another turnstile illustrated with the cyan area, whose backwards map is shown in green, and forwards map in blue.
In order to understand the chaotic transport of field lines through an island chain, we take a zoomed-in look at a single island indicated by the box in figure 2. Figure 3 shows the two period-3 fixed points labeled $\mathbf{x}^{1}_{0}$ and $\mathbf{x}^{2}_{0}$ and the stable and unstable manifolds associated with them. The area around the island o-point, and bounded by the manifolds, is called the island resonance zone (for the exact definition see [55]). Consider the orange area labeled $\Sigma_{\text{inner}}$, on the bottom, defined by the intersections of the inner manifolds. If we apply the forward map $f^{3n}(\Sigma_{\text{inner}})$, it is mapped to the to the red areas, and when we apply the backwards map, it is mapped to the yellow areas.
@@
Locating the heteroclinic orbits can be done by computing the complete manifolds, and locating their intersections. Calculating the manifolds is a relatively computationally costly endeavor, as it requires integrating a large number of trajectories for many iterations. For iterated maps this is not an issue, but if the map is calculated by the numerical integration of trajectories, the computational cost can be significant. Therefore, we implement an algorithm to find these heteroclinic orbits directly with many fewer required integrations.
@@
(here the superscript denotes the fixed point label). We choose $\epsilon_{u}$ small enough to be in the linear regime, here $10^{-6}$, and we then compute $f^{mn_{\text{map},u}}(\mathbf{x}_{u})$, where $m=3$ is the order of the fixed point, and $n_{\text{map},u}$ is the number of times this initial unstable point is mapped back to the same intersection.
@@
and calculate $f^{-m\cdot n_{\text{map},s}}(\mathbf{x}_{s})$.
If $\mathbf{x}_{s}$ and $\mathbf{x}_{u}$ are heteroclinic points, then $f^{m\cdot n_{\text{map},s}}(\mathbf{x}_{u})=f^{-m\cdot n_{\text{map},u}}(\mathbf{x}_{u})$, but this is unlikely to be the case for an initial guess. We therefore minimize the components of:
$\Delta=f^{m\cdot n_{\text{map},u}}(\mathbf{x}_{0}^{1}+\epsilon_{u}\cdot\mathbf{v}_{u}^{1})-f^{-m\cdot n_{\text{map},s}}(\mathbf{x}_{0}^{2}+\epsilon_{s}\cdot\mathbf{v}_{s}^{2}).$ (16)
@@
We choose $n_{\text{map},s}=n_{\text{map},u}=8$ for both the stable and the unstable manifold integration, and the values of $\epsilon_{s,u}$ are initially set to $10^{-6}$. The root-finding converges resulting in a heteroclinic trajectory with 16 points. There are at least two heteroclinic trajectories (one where the stable manifold intersects the unstable in the positive, the other in the negative). To find the other heteroclinic orbit, we perform a new search with $n_{\text{map},s}=8$ for the stable trajectory and $n_{\text{map},u}=7$, and an initial guess for $\epsilon_{s,u}$ halfway between the first and second point of the heteroclinic trajectory that is already found. This converges to a second heteroclinic trajectory. The heteroclinic trajectories thus found are shown in figure 3 as the blue points.
@@ V Chaotic transport in a perturbed tokamak
In this section, we will analyze chaotic transport in a tokamak-like field that has been perturbed. We will first describe the analytical expressions used to calculate an axisymmetric, tokamak-like field and analytical perturbations to it. We calculate the manifolds in this perturbed system and describe an algorithm to find the homoclinic orbits, the intersections of these manifolds. Using the homoclinic orbits, we describe an efficient algorithm to calculate the turnstile area using an action principle by MacKay, Meiss and Percival [33].
In a three-dimensional field, the fast-evaluating function such as (9), is replaced by an expensive integration of the trajectory as in (1). Though the variables $(\psi,\vartheta)$ are replaced with $(R,Z)$, all of the discussed algorithms remain the same. We will start with a axisymmetric field, that resembles the configuration in a tokamak. Calculation of the manifolds in tokamak geometry has been studied theoretically by Punjabi and Boozer [6], Wei and Liang [56], and their signatures have been shown in experiments [19]. Here we build upon this work by implementing our efficient method of calculating the turnstile area, and hence a measure for the degree of chaos and stochastic transport.
@@
which is chosen such as to give a simple expression for the rotational transform $\imath$ or its inverse, $q=1/\imath$ which is more often used in tokamak physics:
@@
Here $K$ and $E$ are the complete elliptic integral of the first and second kind respectively, $\alpha^{2}=(R_{l}-R)^{2}+(Z-Z_{l})^{2}$, $\beta^{2}=(R_{l}+R)^{2}+(Z-Z_{l})^{2}$, and $k=1-\alpha^{2}/\beta^{2}$.
Equations (17) and (21) are implemented numerically in terms of their respective vector potentials, and the magnetic field is calculated through $\mathbf{B}=\nabla\times\mathbf{A}$ using automatic differentiation provided by JAX [58]. We choose $R_{0}=6$, $q_{a}=1.16$, $s=1.2$ for the axisymmetric field, and for the separatrix coil we choose $R_{l}=6$, $Z_{l}=-5.5$, $I=10\pi/\mu_{0}$. The resultant field is integrated numerically, and a Poincaré section is shown in figure 4 (a). The fixed points of the 1-map of this field are also located, using an initial guess found manually, and these are the magnetic axis shown with a blue dot, the location of the separatrix coil shown with a red square, and a hyperbolic x-point in between, shown with an orange cross.
[…]
Figure 4: Poincaré section of the axisymmetric analytical field and the result of applying a perturbation. (a) Axisymmetric field $\mathbf{B}_{\rm axi}$ with the o-point of the main plasma in blue, the separatrix coil location indicated in red, and the x-point it creates identified by an orange cross. The unstable and stable manifolds that enclose the plasma overlap, and are shown by a magenta and blue line respectively. (b) Effect of applying a $m=6$, $n=1$ perturbation field. The direction the perturbation field is indicated by arrows located on the unperturbed separatrix and on a circle in the plasma region. The perturbation creates islands inside the plasma region, and the formation of chaos around the separatrix. The stable and unstable manifolds (magenta and blue respectively) no longer overlap. (c) safety factor $q$ (inverse of rotational transform $\imath$ measured along a horizontal line from the axis, in both the unperturbed (blue) and perturbed field (orange). The islands in the perturbed field correspond with the regions of flat $q$ at integer values of $q$.
Because the fields given by equations (17) and (21) are axisymmetric, the resultant Poincaré section in fig. 4 (a) does not exhibit any chaos, and there are no turnstiles. Now we add a perturbation that depends on the toroidal angle $\phi$ with the form:
@@
where $\phi$ is the usual toroidal angle and $\theta$ is the poloidal angle; $\tan(\theta)=(Z-Z_{0})/(R-R_{0})$ (with $(R_{0},Z_{0})$ the location of the magnetic axis, blue circle in fig. 4 (a)). The radial distribution $\mathfrak{f}(\rho)$ of the perturbation can have any form, we choose one that places the perturbation smoothly everywhere in the configuration by using a Maxwell-Boltzmann probability density function with distribution parameter $d$:
$\mathfrak{f}(\rho,d)=\frac{\sqrt{2}}{\sqrt{\pi}}\frac{\rho^{2}}{d^{3}}e^{\frac{-\rho^{2}}{2d^{2}}}$ (23)
@@
$\mathbf{B}_{\rm pert}$ lies in the $R,Z$-plane, and is tangent to contours of constant $R^{2}A^{\phi}_{\rm pert}$. The perturbation field oscillates $m$ times in one $R,Z$-plane, and this pattern rotates $n$ times as the toroidal angle goes from 0 to $2\pi$. We choose $\phi_{0}=0$, $\theta_{0}=0$, distribution width $d=2$, and mode numbers $m=6$ and $n=1$. This is shown in figure 4 (b).
When this perturbation is applied, the field becomes chaotic, as is seen in figure 4 (b). Islands have opened up on several magnetic surfaces, and the region around the separatrix surface has become chaotic. The x-point is no longer clearly visible in the Poincaré section. Nevertheless, the fixed point is still there, and it is found using the Newton method described in section III.1, and it is indicated with a orange cross.
In section IV we considered a heteroclinic intersection, where stable manifolds of one fixed point intersect with the unstable manifolds of a second point in the same orbit, but here, we can consider the simpler where the stable and unstable manifolds originate from the same point; a homoclinic connection. We find the homoclinic trajectory in this integrated map using the same algorithm as described in IV.1, and calculate the manifolds in the same way as described in section IV. The results of this calculation are shown in figure 4 (b).
@@ V.1 Turnstile area calculation
When the trajectories are calculated by integrating field lines in three dimensions, resolving the stable and unstable manifolds between two homoclinic orbits can be computationally expensive. One would first need to find the homoclinic points (as can be done with the algorithm described in IV.1), and then integrate 50-100 trajectories between them.
[…]
@@
In the case of magnetic fields, the Lagrangian form is can be calculated by integrating the vector potential along the field line:
$\lambda(m_{t})=\int_{\phi=0}^{\phi=nT}\mathbf{A}(\mathbf{x})\cdot\mathrm{d}l.$ (26)
where the integration is along the field line, it is started at $\mathbf{x}=m_{t-1}$ and the vector potential is evaluated on each point of the field line as it is evolved simultaneously with the calculation of eq. (26).
It seems rather remarkable that the computation of a complicated area in phase space can be performed by the simple summation (equation (25)) of action integrals (equation (26)). Though the formulation by MacKay, Meiss, and Percival [33] in terms of Lagrangian forms and area forms holds in any dimension, in the case of a three-dimensional flow and a two-dimensional map, this formula has an intuitive explanation.
Figure 5: Illustration of the calculation of a turnstile area. The area $\Sigma_{n}$ shown in gray is bounded by turnstile sections $\gamma_{n}$ and $\tau_{n}$ between homoclinic points $h_{n}$ and $m_{n}$, is given by the integral of the vector potential over the curve $\tau_{n}\cup-\gamma_{n}$. This integral can be modified to run over $\tau_{n}\cup\zeta_{n}\cup\gamma_{n-1}\cup\xi_{n}-1$. This can be repeated until $\gamma_{n-i}$ is in the linear regime of the fixed point. Similarly the integration over $\tau_{n}$ can be moved forward to $\tau_{n+j}$ in the linear regime of the stable manifold. (bottom): illustration of the turnstile area calculation where the intersecting manifolds have been laid flat.
Figure 5 illustrates how the calculation of the turnstile lobe can be understood. To calculate the flux through the dark gray shaded area, which without loss of generality we label the $n$-th occurrence of the lobe, one can use Stokes’ theorem and integrate the vector potential $\mathbf{A}$ over $\gamma_{n}$, a section of the unstable manifold, and $\tau_{n}$, a section of the stable manifold:
@@
We can now shift the integral over $\gamma_{n}$ by transporting the entire curve backward along the field lines. As this does not change the flux through the area, because the transport is along the field, the result should be identical, and we only need to add the contribution of $\int\mathbf{A}\cdot\mathrm{d}l$ along the endpoints, which move along the trajectories $\zeta_{n-1}$ and $\xi_{n-1}$. We transport the path $\gamma_{n}$ all the way back $2pi$, until it reaches the same section. This integral now becomes:
@@
Instead of integrating the vector potential over two curve segments, whose geometry has to be found by integrating many trajectories, we can instead identify two specific trajectories along which to integrate the vector potential. One issue in the practical application of equation (26) is when to terminate the sum. In practice only about 10 applications of the map suffice to bring a point from the linear regime of the unstable manifold into the linear regime of the stable manifold. Increasing the number of points in the homoclinic trajectories will make the sum converge, but at a large computational cost. Instead, when we are in the linear regime, we can close the sum by evaluating $\int\mathbf{A}\cdot\mathrm{d}l$ along the straight line connecting the homoclinic points in the linear regime near the fixed point.
[…]
@@
Figure 6: The effect of perturbations on the turnstile area using the same axisymmetric field as in figure 4. (a): Each column has a different poloidal mode number $m$, and each row has a different amplitude of the perturbation. Different mode numbers affect the field and chaotic transport differently. Low $m$ mode numbers are resonant with the surfaces inside, and create chaotic fields throughout the volume. (b): Turnstile area vs perturbation amplitude. The response to perturbation is mostly linear, but can act nonlinearly at very high amplitudes. Gaps are caused when the calculation fails. (c): Turnstile area at fixed perturbation amplitude of $\max\left(\frac{{|B_{{\rm pert}}|}}{{|B_{{\rm axi}}|}}\right)=0.004$. For a given perturbation amplitude the $m=4$ perturbation creates the largest $\Sigma_{\text{turnstile}}$, followed by the $m=6$ mode.
@@
We now use this algorithm to compare the effects of varying the perturbation mode numbers and perturbation amplitude on the turnstile area in the tokamak toy model equilibrium. To compare different perturbations, we normalize the perturbation strength the axisymmetric field strength, by evaluating the maximum of $|B_{\rm pert}|/|B_{\rm axi}|$ on the unperturbed separatrix. Poincaré sections at three amplitudes, for the mode numbers $m=2,4,7,9$ ($n$ always 1) are shown in figure 6 (a). All perturbations create chaotic regions, but the perturbations with lower $m$ create more chaos in the core. The rotational transform crosses the rational values $m/1$ for those $m<7$ and the mode couples with resonant surfaces there.
@@
The size of the turnstile area as a function of the mode number is shown in figure 6 (b), The area scales approximately linearly with the applied perturbation, though some modes show nonlinearity at high amplitudes. Gaps in the graphs are caused by the failure of the algorithm, which occurs sporadically, and is observed to happen when the manifolds are close to intersecting in additional locations, as can be seen close to happening in the bottom row of panel a). Despite these few failures, the method is able to calculate turnstile areas over a large range of perturbation amplitudes, and at the maximum applied perturbations the perturbation field is is around 15% of the unperturbed field strength. We thus have a reliable method to evaluate the turnstile area over a broad range of perturbation amplitudes, from nearly integrable fields, to moderately chaotic fields up to almost completely chaotic fields.
@@
Different perturbations are more or less effective at creating chaotic transport on the separatrix. To compare this, we plot the turnstile area at a fixed perturbation amplitude of $\max\left(\tfrac{|B_{\rm pert}|}{|B_{\rm axi}|}\right)=0.0040$ for all different modes. This is shown in figure 6 (c).
@@
Understanding which modes are most effective at creating chaos in fields is an important and long-standing area of research in tokamak physics, where so-called Resonant Magnetic Perturbation (RMP) [15] coils are placed around the vessel to create a chaotic edge. Computing the lobe structures is key to predicting the heat footprints on divertor targets [59]. We expect that the turnstile area calculation could be an effective tool in designing RMP fields. Furthermore, given the potential risk posed by runaway electrons for future tokamak experiments, there is a big effort in studying mitigation methods, and one proposal are Passive Runaway Electron Mitigation Coil (PREMCs) [60]. Here a shaped coil is placed near the vessel, and the current induced in it by the abrupt changes in flux brought on by the disruptive events preceding the formation of a dangerous beam of near-relativistic electrons, will induce a current in the coil which creates chaos in the field, disrupting the beam. Such a coil could be optimized to create maximum chaotic field at desired locations.